According to government data, 22% of American children under the age of 6 live in households with incomes less than the official poverty level. A random sample of 15 children is selected for a study of learning in early childhood. Assuming the distribution is binominal, calculate the probability that exactly 3 children in the sample come from poverty-level households.
Answer #1

Solution :

Given that n = 15, p = 0.22

=> q = 1 - p = 0.78

=> From binomial distribution, P(X = r) = nCr*p^r*q^(n-r)

==> P(X = 3) = 15C3*0.22^3*0.78^12

= 0.2457